OpenStudy (anonymous):

# Solve P = 2(l + w) for l. What are the missing values in the table? (1 point) P w l 14 2 ? 22 8 ?

OpenStudy (anonymous):

(i) p=2(l + w) 14=2l + 2w 14=2l + 2*2 14=2l + 4 10= 2l 10/2=l 5=l (ii) similarly l=3